OpenStudy (vortish):

factor the Binomial completly 144x^2-1

OpenStudy (anonymous):

144x^2-1 = (12x)^2-1 Use a^2 - b^2 = (a+b)(a-b) to factorize it.

OpenStudy (anonymous):

144x^2-1 = (12x)^2-1 =(12x+1)(12x-1)